# MA-LoT: Multi-Agent Lean-based Long Chain-of-Thought Reasoning enhances Formal Theorem Proving

### v2

@@ Appendix E Training examples
This section we present the training examples for SFT data and correction data. We can see that in the LoT-TL training stage, we keep the system prompt to switch off the Long CoT thinking. In particular, in the SFT data, we switch off the entire Long CoT, in the correction data training, we only switch off the analysis of error message in the Long CoT while provide the error message as well as the original proof draft as a component in the Long CoT.
E.1 LoT-TL Training example for SFT data
@@
E.2 LoT-TL Training example for correcton data
@@
Appendix F Agents input example
F.1 Prover agent input-output example
@@
F.2 Corrector agent input-output example

### v3

@@ Appendix F Training examples
This section, we present the training examples for SFT data and correction data. We can see that in the LoT-TL training stage, we keep the system prompt to switch off the Long CoT thinking. In particular, in the SFT data, we switch off the entire Long CoT., In the correction data training, we only switch off the analysis of error messages in the Long CoT while providing the error message as well as the original proof draft as a component in the Long CoT.
F.1 LoT-TL Training example for SFT data
@@
F.2 LoT-TL Training example for correction data
@@
Appendix G Models input example
G.1 Prover model input-output example
@@
G.2 Corrector model input-output example